OpenStudy (anonymous):

what are the dimensions of the largest rectangular field that can be enclosed with 60m of wire?

OpenStudy (anonymous):

let length be x and breadth be y ... Perimeter =60 m 2(x+y)=60 x=30-y ...i Now, Area of a rectangle =x*y A=(30-y)*y A=30y-y^2

OpenStudy (anonymous):

Now, u need to find the value of y for which A is maximum. can u?

OpenStudy (anonymous):

@vianca22 Did u get up to what I did?